Venture Capital:
Henry invested $12000 in a new resturant. When the resturant was sold two years later, he received $27000. Find the average annual return by solving the equation 12,000(1+r)^2=27000.
used the foil to solve the (1+r)^2 or (1+r)(1+r)
12000(1+2r+r^2)=27000
12000+24,000r+12000r^2=27000
-15000+24,000r+1200r^2=0
THis is as far as I could get. Not sure my next step.

Answer by scott8148(6628)

dividing by 12000 ___ (1 + r)^2 = 9 / 4

taking sqrt ___ 1 + r = 3 / 2

subtracting 1 ___ r = 1 / 2 ___ or .5 ___ or 50%

pretty good investment...